Our visitors are already leaving tomorrow – University of Copenhagen

Forward this page to a friend Resize Print Bookmark and Share

EastGrip - The East Greenland Ice-core Project > Field Diaries Folder > UK diaries 2016 > Our visitors are alrea...

24 May 2016

Our visitors are already leaving tomorrow

Tuesday 24th May 2016

Malte, Joel, Chris and Bjarni (Twin Otter pilot) enjoy waiting for the snow lamb to be ready.

The PARCA team have been very efficient. Today, they flew to one site to maintain a station and they were back early afternoon. Chris with the Pistenbully and crane helped to move the station from 4 km away to a point 300 m from EGRIP camp. This station move took only a few hours.

Although everybody was busy, cutting snow, making shelves, maintaining pumps and blowing snow, Joel still found time to prepare a snow lamb for dinner. Some tin foil is laid out in the snow, then charcoal is added and lit. A spiced lamb leg is wrapped in foil and added to the charcoal. Cooking time is about 1 meter, i.e. the time it takes for the charcoal with lamb on top to melt the snow to one meter depth. Dinner was a nice affair and we enjoyed the company of our visitors. They are already leaving tomorrow and will be heading to Qaanaaq after a brief visit to maintain the station at the old NEEM site.

What we have done today:

  1. Assisted Twin Otter and crew. They received a total of 2,200 liter fuel.
  2. Helped move PARCA weather station close to EGRIP camp.
  3. Made walls straight and installed shelves in cooks freezer.
  4. Excavated German pit to 4 m depth.
  5. Maintaining pumps.
  6. Making a level base around staircase shaft.
  7. Exchanged oil pressure sensor on main generator.

Ad.7: Already a few days after arrival our main generator began to make spontaneous stops, due to a low oil pressure alarm. We soon located the source as a defect oil pressure sensor. We exchanged it with a spare, but the show continued, quite often the generator stalled in middle of the night. By contacting the company, it turned out that both our sensors came from a faulty batch. Yesterday, the PARCA team brought us two new sensors, and we decided to change sensors during a controlled stop after lunch. However, Chris and J.P. were in the generator room as the generator again decided to stop, and the sensor was exchanged immediately. Since then, no problems.

Weather today: Mostly clear with few high clouds. Temp. –16 °C to -26 °C, Wind 5kt WSW to 15 kt from SSW. Visibility: To horizon.

FL, J.P. Steffensen

← Previous entry   Next entry →